OpenStudy (kamiya):

The ratio of the perimeter of two similar squares is 5 to 4. If the area of the smaller square is 32 square units, what is the area of the larger square?

OpenStudy (anonymous):

It's 50 sq units.

OpenStudy (asnaseer):

The ration of areas is equal to the square of the ratios of lengths, so you can use:\[\frac{5^2}{4^2}=\frac{x}{32}\]and solve for x. @Mertsj I /believe/ you got the right-hand-side fraction inverted?

OpenStudy (asnaseer):

*ratio NOT ration

OpenStudy (mertsj):

Oh, yes, I did. Thanks asnaseer

OpenStudy (asnaseer):

np